OpenStudy (erniellyd28):

if a sample of a certain gas is heated.it will expand.the formula v(t)=0.175T+50 calculates the volume in cubic inches of a sample of gas temperture T in degrees celsuis. a)evaluate v(o) and interpret the result. b)if the temperature increases by 10 degrees celsuis. by how much does the volume increase? c) what is the volume of the gas when the temperature is at 100 degrees celsius?

OpenStudy (peachpi):

v(0) is asking for the initial volume. What do you get if you plug in 0 for T?

OpenStudy (erniellyd28):

0?

OpenStudy (peachpi):

no, what does the right side of the equation simplify to? v(0)=0.175(0)+50

OpenStudy (erniellyd28):

50?

OpenStudy (peachpi):

yes

OpenStudy (peachpi):

so the initial volume is 50 cubic inches

OpenStudy (peachpi):

sorry that's not the initial volume. It's the volume at 0 degrees celsius

OpenStudy (peachpi):

for B) there are 2 steps. First plug in 10 for T to get the volume at 10 °C

OpenStudy (erniellyd28):

im not sure how to do that i'm very sorry

OpenStudy (peachpi):

Start with your formula v(t)=0.175T+50 Replace T with 10 v(10)=0.175(10)+50 Now what do you get for the right side?

OpenStudy (erniellyd28):

51.75?

OpenStudy (peachpi):

yes. the next step is to subtract the 50 you got from part A, since they want the increase. So do 51.75 - 50

OpenStudy (erniellyd28):

1.75

OpenStudy (peachpi):

right

OpenStudy (peachpi):

For C, plug in 100 for T T(100) = 0.175(100) + 50

OpenStudy (erniellyd28):

67.5?

OpenStudy (peachpi):

yes

OpenStudy (erniellyd28):

so the answer is 67.5?

OpenStudy (peachpi):

yes

OpenStudy (erniellyd28):

if i could give you more medals i would :-)

OpenStudy (erniellyd28):

thank you very much

OpenStudy (peachpi):

:) you're welcome

OpenStudy (erniellyd28):

the computer said the answer was wrong

OpenStudy (erniellyd28):

it said it was 50 not 67.5